OpenStudy (anonymous):

MEDAL A graduated cylinder with a radius of 3 cm contains water to a height of 9 cm. Yann adds 15 steel balls, each with a diameter of 2.54 cm. What is the total volume of the contents in the cylinder after the balls are added? Use 3.14 to approximate pi and round your answer to one decimal place. cm3

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

Wait lemme c if I can solve this :P

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

Sorry I can't D: I apologize.... @directrix can u help

Directrix (directrix):

Let me see.

Directrix (directrix):

Take the equation for the cylinder, r = 3 and h = 9 Get the volume of one sphere, r = 1.27. There are 15 spheres.

Directrix (directrix):

That gives this: pi * 3^2 * 9 + 15 * ( (4/3) * 3.14 * (1.27)^3 ) = ?

OpenStudy (anonymous):

???

Directrix (directrix):

Cylinder volume: pi * 3^2 * 9 One Sphere: ( (4/3) * 3.14 * (1.27)^3 )

Directrix (directrix):

15 spheres: 15 * ( (4/3) * 3.14 * (1.27)^3 )

Directrix (directrix):

@TeenWolfGirl Evaluate this: 3.14 * 3^2 * 9 + 15 * ( (4/3) * 3.14 * (1.27)^3 ) = ?

OpenStudy (anonymous):

383?

Directrix (directrix):

@Godlovesme Will you evaluate: 3.14 * 3^2 * 9 + 15 * ( (4/3) * 3.14 * (1.27)^3 ) =

Directrix (directrix):

Note that I subbed 3.14 for pi prior to the calculation.

OpenStudy (anonymous):

383

OpenStudy (anonymous):

wait

Directrix (directrix):

You can take the expression to the "Wolf" and calculate there: http://www.wolframalpha.com/

Directrix (directrix):

>> round your answer to one decimal place

OpenStudy (anonymous):

382

OpenStudy (anonymous):

i did look at the website link that is where it is from

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

382.9784524 =383 right?

Directrix (directrix):

Yes. To the nearest tenth, that is what?

Directrix (directrix):

383.0 to the nearest tenth if you approximate pi as 3.14 before number crunching. @TeenWolfGirl

OpenStudy (anonymous):

Thanks

Directrix (directrix):

You are welcome.